OpenStudy (anonymous):

Possibly Mathematica specific. Here f[x] is a mathematica least squares fit to a dataset. "Use f[x]=E^(5 + .5 x)​ to get a good fit with a solution of the logistic differential equation" GIVEN Solution to the logistic differential equation y[x] -> (b*E^(r*x + b*k))/(-1 + E^(r*x + b*k))​ Now somehow, I am supposed to reconcile F[x] into y[x], and create some kind of logistic model. Is it just a case of identifying the relative variables between the equations, and then use substitution, or is there more to this, requiring some manipulation?

OpenStudy (nincompoop):

beyond my mathematicabilities

OpenStudy (nincompoop):

@Miracrown

Miracrown (miracrown):

Do you have an a dataset to work from? Or do you need need to come up with a mathematica command line?

OpenStudy (anonymous):

I ahve a dataset

OpenStudy (anonymous):

Should I show the full data and conversion that led up to the creation of an f[x] =?

Miracrown (miracrown):

no I don't think that would be helpful you want to take the assumption that your solution function is log based they even say "solution to the logistic differential equation"

Miracrown (miracrown):

Right you can practice on a simpler function, then once you work through it you can use the data set you have

OpenStudy (anonymous):

probably better to keep with the simplified version ?

OpenStudy (anonymous):

cool

Miracrown (miracrown):

Right they told mathematic to fit a function to the data set that is the line fitter[t_] = E^Fit[logdata,{1,t},t]

OpenStudy (anonymous):

wow, you're actually going to answer this.. this is awesome.

Miracrown (miracrown):

LOL

OpenStudy (anonymous):

or at least get me clued in.

Miracrown (miracrown):

this looks like it is the solution

OpenStudy (anonymous):

so mathematica did it already? the fit command actually did a solution to the logistic differential equation.

Miracrown (miracrown):

so this is a semi log plot it should give you a straight line since it is a population time graph so the function should look like f(t) = A e^rt where r is the slope of the semi log plot t is in years and A is the y intercept of the plot

Miracrown (miracrown):

\(\color{blue}{\text{Originally Posted by}}\) @hughfuve so mathematica did it already? the fit command actually did a solution to the logistic differential equation. \(\color{blue}{\text{End of Quote}}\) It should have, it should give you either an equation or the values of A and r

Miracrown (miracrown):

I think it should kick out an equation from what is here

OpenStudy (anonymous):

okay hmmm.. so fitter[t_] = E^Fit[logdata,{1,t},t] gave me E^(4.382968909053888 + 0.01300265909581666*x)​ and that should probably be A e^rt where r is the slope of the semi log plot t is in years and A is the y intercept of the plot

OpenStudy (anonymous):

so that constant on the front might be A E^(4.382968909053888) E^(0.01300265909581666*x)​

Miracrown (miracrown):

well A = e^4.38 and R = .013

OpenStudy (anonymous):

sweet.. concur

Miracrown (miracrown):

not quite we raise e to the 4.38 to get A then r is .0130026

OpenStudy (anonymous):

gotcha.. we are on the same page there then.

Miracrown (miracrown):

Mhm

OpenStudy (anonymous):

so as for this equation.. y[x] -> (b*E^(r*x + b*k))/(-1 + E^(r*x + b*k))​ we are basically already in a simplified version of this?

Miracrown (miracrown):

right we get the simplified form

OpenStudy (anonymous):

or rather, we are now in a simplified version of this.

Miracrown (miracrown):

it looks like the way they have it A = b but r is the same and I think k is 0

Miracrown (miracrown):

and x is t

OpenStudy (anonymous):

gotcha

Miracrown (miracrown):

k might be something very very small it might be 1 x 10^-12 depending on how far mathematica calculates it out

OpenStudy (anonymous):

a limit of it's precision you think?

OpenStudy (anonymous):

well miracrown.. that has put me in a much clearer vision.. thank you.

Miracrown (miracrown):

right k should go to 0 for a perfect fit then all the b*k terms go away :)

Miracrown (miracrown):

and our equation becomes b*e^rt or b* e^rx

OpenStudy (anonymous):

awesome.. thank you.. thank you.

Miracrown (miracrown):

Yw yw :)